# Burger[Uncaptioned image]: Robust Graph Denoising-augmentation Fusion and Multi-semantic Modeling in Social Recommendation

### v2

@@ 3.5. Graph Denoising-augmentation Fusion
(24) $\forall v^{-}\in\mathcal{P}_{u},w\in(\mathcal{C}_{u}-\mathcal{P}_{u}),pf_{u}(v^{-})\geq pf_{u}(w),$
where $|\mathcal{P}_{u}|=|\mathcal{B}_{u}|$, and $\mathcal{C}_{u}-\mathcal{P}_{u}$ represents the set of potential non-friends of user $u$. To simultaneously mine friends from $\mathcal{P}_{u}$ who have shared item preferences with $u$ and remove friends from $\mathcal{B}_{u}$ who don’t have shared item preferences with $u$, first combine $\mathcal{P}_{u}$ with $\mathcal{B}_{u}$ into one set and select the $|\mathcal{B}_{u}|$ users with the highest $\phi_{uv^{-}}$ from this set. The selected users are denoted as the set ^$\hat{\mathcal{B}}_{u}$. It is evident that set ^$\hat{\mathcal{B}}_{u}$ satisfies the following constraints:
(25) $\forall v^{-}\in\hat{\mathcal{B}}_{u},w\in(\mathcal{B}_{u}\cup\mathcal{P}_{u}-\hat{\mathcal{B}}_{u}),\phi_{uv^{-}}\geq\phi_{uw},$
where |^$|\hat{\mathcal{B}}_{u}|=|\mathcal{B}_{u}|$. Repeat the above procedure to calculate ^$\hat{\mathcal{B}}_{u}$ for each user in the social graph, according to Eq.23, Eq.24 and Eq.25. Then, based on all users’ ^$\hat{\mathcal{B}}_{u}$, a new denoising-augmentation fusion social graph is constructed.
4. Experiments
@@
4.1. Experimental Settings
4.1.1. Datasets.
Extensive experiments are conducted on three real datasets: Ciao, Douban, and Yelp. The data is split into training, test sets with a ratio of 7:3. The statistics of the three datasets are shown in Table 1.
Table 1. Table 1: Statistics information of datasets.
@@ 4.1. Experimental Settings
4.1.2. Evaluation Metrics
To assess the performance of Burger model and other baselines, we utilize commonly used metrics, i.e., Hit Ratio (HR@N) and Normalized Discounted Cumulative Gain (NDCG@N). Following (Qin et al., 2020)’s suggestion, a leave-one-out strategy is adopted to compute the metrics, i.e., only one sample is retained for the target user in the test set. Therefore, based on this strategy, it is sufficient for experiments to select N = 1 and N = 3.
4.1.3. Parameter Settings.
In our experiments, for fairness, in the baseline comparison experiments, the embedding dimensions of all models are set to 512, batch size to 1024, regularization parameter $\lambda_{2}$ to 0.00001, and early stopping patience to 5. The Adam optimizer is used to train the models, while other baseline model parameters are set according to the optimal parameters from the original research. For Burger, the initial social network random enhancement probability is set to 0.01, and the range of convolutional layers for the two graphs is $\left\{1,2,3,4\right\}$. For loss $\mathcal{L}_{\Omega_{1}}$ and $\mathcal{L}_{\Omega_{2}}$, set the constants $C_{1}$ and $C_{2}$ to 1. For more details, please refer to section 4.2.3.
4.1.4. Baselines.
@@ 4.1.2. Evaluation Metrics
  • PMF (Salakhutdinov and Mnih, 2007): It is a probabilistic factorization model that uses matrix factorization to map users and items to a common embedding space.
  • TrustMF (Yang et al., 2017): It decomposes the social and user-item interaction matrices to map users to two separate latent spaces, enhancing recommendation performance.
  • EATNN (Chen et al., 2019): It’s a transfer learning model with adaptive features and an attention network to analyze dynamic user-item interactions.
  • DiffNet (Wu et al., 2019): This model enhances user embeddings iteratively through an influence diffusion mechanism and captures collaborative signals between users and items via a fusion layer.
  • DGRec (Song et al., 2019): The model aims to capture evolving user behaviors and social influences by merging a recurrent neural network with a graph attention layer.
  • NGCF+ (Wang et al., 2019): This method enhances collaborative filtering with GNNs by conducting message propagation across a socially-aware user-item relation graph.
  • KCGN (Huang et al., 2021): Enhancing social recommendations involves merging item knowledge with social influence using a multi-task learning approach.
  • MHCN (Yu et al., 2021): It conducts hypergraph convolution by mining Triangle motifs in the heterogeneous graph.
  • SMIN (Long et al., 2021): This model boosts social recommendations by employing metapath-guided heterogeneous graph learning and self-supervised signals via mutual information maximization.
  • DSL (Wang et al., 2023): It aligns social information with user-item interaction information adaptively by sampling user node pairs.
Table 2. Table 2: Performance comparison of three metrics HR@1, HR@3 and NDCG@3 on three datasets for all methods. * indicates the p-values of t-tests between Burger and the best baselines are smaller than 0.05.
@@
4.2. Recommendation Performance
4.2.1. Overall Performance(RQ1)
Table 2 presents the results of all baselines compared to Burger. The bold-faced values indicate the best results on three datasets, and the underlined value represents the best result among all baselines. From these results, the following conclusions are drawn:
@@ 4.2.1. Overall Performance(RQ1)
(2) Baselines based on self-supervised signals outperform GNNs-based baselines by an average of 23.76%. This indicates that self-supervised signals effectively integrate social networks into user’s item preference modeling, whereas naively characterizing the entire social network leads to suboptimal performance. And GNNs-based baselines outperform matrix factorization baselines by an average of 35.95%, as GNNs capture higher-order, complex nonlinear relations, better fitting user preferences and item characteristics.
4.2.2. Ablation Study(RQ2)
To verify the effectiveness of each module in Burger, we conducted ablation studies. Specifically, we compared Burger with the following variants: w/o s represents the removal of the sliding window based social tensor construction, with a single social graph serving as the model input; w/o $\mathcal{L}_{\Omega_{1}}$ represents the removal of coordinating relative positions in the semantic space $\Omega_{1}$; w/o $\mathcal{L}_{\Omega_{2}}$ represents the removal of coordinating relative positions in the semantic space $\Omega_{2}$; w/o g represents the removal of graph denoising -augmentation fusion. According to these results from Table 3, we can draw several conclusions as follows:
@@ 4.2.2. Ablation Study(RQ2)
Table 3. Table 3: Abaltion studies for different components in Burger.
@@
4.2.3. Hyper-parameters Sensitivity(RQ3)
We explored how the key hyperparameters in the Burger model affect the model’s performance. including $\alpha$, $\beta$, and sliding window size $\tau$. For brevity, we only present results for $\tau$ on the Douban dataset.
@@ 4.2.3. Hyper-parameters Sensitivity(RQ3)
Figure 3. Sensitivity of hyperparameter $\alpha$, $\beta$ and sliding window size $\tau$.
@@
4.3. Further Analysis
4.3.1. Bayesian method(RQ4).
@@ 4.3. Further Analysis
(26) $P(\varphi|\mathfrak{f})$ $=2F(\varphi)f(\varphi),$
(27) $P(\varphi|\mathfrak{nf})$ $=2f(\varphi)[1-F(\varphi)],$
where $P(\varphi|\mathfrak{f})$ and $P(\varphi|\mathfrak{nf})$ are the theoretical distributions of potential friends and potential non-friends, respectively. The detailed derivation process of the two conditional densities is shown in BNS (Liu and Wang, 2023) , which we do not elaborate here due to space limit. Given the set $\mathcal{C}_{u}$ and user $v^{-}\in\mathcal{C}_{u}$, the posterior probability $pf_{u}(v^{-})$ can be calculated as: $pf_{u}(v^{-})=P(\mathfrak{f}|\varphi_{uv^{-}})\propto P(\varphi_{uv^{-}}|\mathfrak{f})P_{\mathfrak{f}}(v^{-})$. Based on Eq.26 and Eq.27, the Bayesian posterior probability can be naturally derived as follows:
$pf_{u}(v^{-})$ $=\frac{P(\mathfrak{f}|\varphi_{uv^{-}})}{P(\mathfrak{f}|\varphi_{uv^{-}})+P(\mathfrak{nf}|\varphi_{uv^{-}})}$
@@
$=\frac{2F(\varphi_{uv^{-}})f(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})}{2F(\varphi_{uv^{-}})f(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})+2f(\varphi_{uv^{-}})[1-F(\varphi_{uv^{-}})]P_{\mathfrak{nf}}(v^{-})}$
@@
Figure 4. Theoretical distribution and actual distribution on three datasets.
4.3.2. The Reliability of Enhanced Social Graph(RQ5).
@@
Figure 5. The reliability verification of enhanced social graph.
It’s crucial to verify if the enhanced social graph has less noise and more reliable latent relations. We input the Burger-enhanced social graph into Diffnet (Wu et al., 2019) and MHCN (Yu et al., 2021), using GDMSR (Quan et al., 2023) as a baseline. According to the Fig.5, we draw the following conclusions: (1) The performance of the Burger-enhanced social graph is superior to that of the original social graph. This can be attributed to the fact that the graph denoising-augmentation fusion module can effectively explore latent social relations and reduce social noise. (2) Compared to the baseline, the superiority of our method lies in the fact that mining potential social relations can alleviate the problem of overly sparse social graph after enhancement. The social graph enhanced by the baseline can lead to insufficient learning of user social representations, and our method effectively solves this problem.
4.3.3. Robustness Evaluation(RQ6).
To evaluate the robustness of Burger, we randomly added a portion of social relations as noise perturbations in the original social network. The synthetic social graph is combined with user-item interaction information to test the performance of Burger. As shown in Table 4, compared with the two baselines, Burger’s robust performance on the synthetic data lies between them, which also indicates that the robustness of Burger is at a relatively considerable level. This can be attributed to the fact that although the synthetic dataset contains a large amount of noise, the graph denoising-augmentation fusion module can still capture this noise.
Table 4. Table 4: Comparison of robustness performance of Burger and other baselines.

### v3

@@ III-E Graph Denoising-augmentation Fusion
$\forall v^{-}\in\mathcal{P}_{u},w\in(\mathcal{C}_{u}-\mathcal{P}_{u}),pf_{u}(v^{-})\geq pf_{u}(w),$ (24)
where $|\mathcal{P}_{u}|=|\mathcal{B}_{u}|$, and $\mathcal{C}_{u}-\mathcal{P}_{u}$ represents the set of potential non-friends of user $u$. To simultaneously mine friends from $\mathcal{P}_{u}$ who have shared item preferences with $u$ and remove friends from $\mathcal{B}_{u}$ who don’t have shared item preferences with $u$, first combine $\mathcal{P}_{u}$ with $\mathcal{B}_{u}$ into one set and select the $|\mathcal{B}_{u}|$ users with the highest $\phi_{uv^{-}}$ from this set. The selected users are denoted as the set $\hat{\mathcal{B}}_{u}$. It is evident that set $\hat{\mathcal{B}}_{u}$ satisfies the following constraints:
$\forall v^{-}\in\hat{\mathcal{B}}_{u},w\in(\mathcal{B}_{u}\cup\mathcal{P}_{u}-\hat{\mathcal{B}}_{u}),\phi_{uv^{-}}\geq\phi_{uw},$ (25)
where $|\hat{\mathcal{B}}_{u}|=|\mathcal{B}_{u}|$. Repeat the above procedure to calculate $\hat{\mathcal{B}}_{u}$ for each user in the social graph, according to Eq.23, Eq.24 and Eq.25. Then, based on all users’ $\hat{\mathcal{B}}_{u}$, a new denoising-augmentation fusion social graph is constructed.
IV Experiments
@@
IV-A Experimental Settings
IV-A1 Datasets
Extensive experiments are conducted on three real datasets: Ciao, Douban, and Yelp. The data is split into training, test sets with a ratio of 7:3. The statistics of the three datasets are shown in Table I.
TABLE I: Table I: Statistics information of datasets.
@@ IV-A1 Datasets
IV-A2 Evaluation Metrics
To assess the performance of Burger model and other baselines, we utilize commonly used metrics, i.e., Hit Ratio (HR@N) and Normalized Discounted Cumulative Gain (NDCG@N). Following [37]’s suggestion, a leave-one-out strategy is adopted to compute the metrics, i.e., only one sample is retained for the target user in the test set. Therefore, based on this strategy, it is sufficient for experiments to select N = 1 and N = 3.
IV-A3 Parameter Settings
In our experiments, for fairness, in the baseline comparison experiments, the embedding dimensions of all models are set to 512, batch size to 1024, regularization parameter $\lambda_{2}$ to 0.00001, and early stopping patience to 5. The Adam optimizer is used to train the models, while other baseline model parameters are set according to the optimal parameters from the original research. For Burger, the initial social network random enhancement probability is set to 0.01, and the range of convolutional layers for the two graphs is $\left\{1,2,3,4\right\}$. For loss $\mathcal{L}_{\Omega_{1}}$ and $\mathcal{L}_{\Omega_{2}}$, set the constants $C_{1}$ and $C_{2}$ to 1. For more details, please refer to section IV-B3.
IV-A4 Baselines
@@
  • PMF [7]: It is a probabilistic factorization model that uses matrix factorization to map users and items to a common embedding space.
  • TrustMF [4]: It decomposes the social and user-item interaction matrices to map users to two separate latent spaces, enhancing recommendation performance.
  • EATNN [38]: It’s a transfer learning model with adaptive features and an attention network to analyze dynamic user-item interactions.
  • DiffNet [22]: This model enhances user embeddings iteratively through an influence diffusion mechanism and captures collaborative signals between users and items via a fusion layer.
  • DGRec [39]: The model aims to capture evolving user behaviors and social influences by merging a recurrent neural network with a graph attention layer.
  • NGCF+ [40]: This method enhances collaborative filtering with GNNs by conducting message propagation across a socially-aware user-item relation graph.
  • KCGN [41]: Enhancing social recommendations involves merging item knowledge with social influence using a multi-task learning approach.
  • MHCN [42]: It conducts hypergraph convolution by mining Triangle motifs in the heterogeneous graph.
  • SMIN [43]: This model boosts social recommendations by employing metapath-guided heterogeneous graph learning and self-supervised signals via mutual information maximization.
  • DSL [20]: It aligns social information with user-item interaction information adaptively by sampling user node pairs.
TABLE II: Table II: Performance comparison of three metrics HR@1, HR@3 and NDCG@3 on three datasets for all methods. * indicates the p-values of t-tests between Burger and the best baselines are smaller than 0.05.
@@
IV-B Recommendation Performance
IV-B1 Overall Performance(RQ1)
Table II presents the results of all baselines compared to Burger. The bold-faced values indicate the best results on three datasets, and the underlined value represents the best result among all baselines. From these results, the following conclusions are drawn:
@@ IV-B1 Overall Performance(RQ1)
(2) Baselines based on self-supervised signals outperform GNNs-based baselines by an average of 23.76%. This indicates that self-supervised signals effectively integrate social networks into user’s item preference modeling, whereas naively characterizing the entire social network leads to poor performance. And GNNs-based baselines outperform matrix factorization baselines by an average of 35.95%, as GNNs capture higher-order, complex nonlinear relations, better fitting user preferences and item characteristics.
IV-B2 Ablation Study(RQ2)
To verify the effectiveness of each module in Burger, we conducted ablation studies. Specifically, we compared Burger with the following variants: w/o s represents the removal of the sliding window based social tensor construction, with a single social graph serving as the model input; w/o $\mathcal{L}_{\Omega_{1}}$ represents the removal of coordinating relative positions in the semantic space $\Omega_{1}$; w/o $\mathcal{L}_{\Omega_{2}}$ represents the removal of coordinating relative positions in the semantic space $\Omega_{2}$; w/o g represents the removal of graph denoising -augmentation fusion. According to these results from Table III, we can draw several conclusions as follows:
@@ IV-B2 Ablation Study(RQ2)
TABLE III: Table III: Abaltion studies for different components in Burger.
@@
IV-B3 Hyper-parameters Sensitivity(RQ3)
We explored how the key hyperparameters in the Burger model affect the model’s performance. including $\alpha$, $\beta$, and sliding window size $\tau$. For brevity, we only present results for $\tau$ on the Douban.
@@ IV-B3 Hyper-parameters Sensitivity(RQ3)
Figure 3: Sensitivity of hyperparameter $\alpha$, $\beta$ and sliding window size $\tau$.
@@
IV-C Further Analysis
IV-C1 Bayesian method(RQ4)
@@
$P(\varphi|\mathfrak{f})$ $=2F(\varphi)f(\varphi),$ (26)
$P(\varphi|\mathfrak{nf})$ $=2f(\varphi)[1-F(\varphi)],$ (27)
where $P(\varphi|\mathfrak{f})$ and $P(\varphi|\mathfrak{nf})$ are the theoretical distributions of potential friends and potential non-friends, respectively. Given the set $\mathcal{C}_{u}$ and user $v^{-}\in\mathcal{C}_{u}$, the posterior probability $pf_{u}(v^{-})$ can be calculated as: $pf_{u}(v^{-})=P(\mathfrak{f}|\varphi_{uv^{-}})\propto P(\varphi_{uv^{-}}|\mathfrak{f})P_{\mathfrak{f}}(v^{-})$. Based on Eq.26 and Eq.27, the Bayesian posterior probability can be naturally derived as follows:
$pf_{u}(v^{-})$ $=\frac{P(\mathfrak{f}|\varphi_{uv^{-}})}{1}$
$=\frac{P(\mathfrak{f}|\varphi_{uv^{-}})}{P(\mathfrak{f}|\varphi_{uv^{-}})+P(\mathfrak{nf}|\varphi_{uv^{-}})}$
@@
$=\frac{F(\varphi_{uv^{-}})f(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})}{F(\varphi_{uv^{-}})f(\varphi_{uv^{-}})P_{\mathfrak{f}}(v^{-})+[1-F(\varphi_{uv^{-}})]f(\varphi_{uv^{-}})P_{\mathfrak{nf}}(v^{-})}$
@@
Figure 4: Theoretical distribution and actual distribution on three datasets.
IV-C2 The Reliability of Enhanced Social Graph(RQ5)
@@
Figure 5: The reliability verification of enhanced social graph.
It’s crucial to verify if the enhanced social graph has less noise and more reliable latent relations. We input the Burger-enhanced social graph into Diffnet [22] and MHCN [42], using GDMSR [18] as a baseline. According to the Fig.5, we draw the following conclusions: (1) The performance of the Burger-enhanced social graph is superior to that of the original social graph. This can be attributed to the fact that the graph denoising-augmentation fusion module can effectively explore latent social relations and reduce social noise. (2) Compared to the baseline, the superiority of our method lies in the fact that mining potential social relations can alleviate the problem of overly sparse social graph after enhancement. The social graph enhanced by the baseline can lead to insufficient learning of user social representations, and our method effectively solves this problem.
IV-C3 Robustness Evaluation(RQ6)
To evaluate the robustness of Burger, we randomly added a portion of social relations as noise perturbations in the original social network. The synthetic social graph is combined with user-item interaction information to test the performance of Burger. As shown in Table IV, compared with the two baselines, Burger’s robust performance on the synthetic data lies between them, which also indicates that the robustness of Burger is at a relatively considerable level. This can be attributed to the fact that although the synthetic dataset contains a large amount of noise, the graph denoising-augmentation fusion module can still capture this noise.
TABLE IV: Table IV: Comparison of robustness performance of Burger and other baselines.
@@ V Conclusion
Acknowledgement
This work is supported by Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.